OpenStudy (anonymous):

FAN AND MEDAL 0.5% as a fraction I think it's 1/20 or 1/500

OpenStudy (anonymous):

The fraction is 1/200.

OpenStudy (anonymous):

Simplify the fraction by multiplying numerator and denominator by 2 and you get 1/200.

OpenStudy (anonymous):

thank you @StudyBuddy17

OpenStudy (er.mohd.amir):

0.5%=0.5/100=5/1000=1/200

OpenStudy (anonymous):

1/200 is your answer

OpenStudy (anonymous):

https://mathway.com/

OpenStudy (anonymous):

Your welcome alibaby.